6. Goode Inc.’s stock has a required rate of return of 11.50%, and it sells for $25.00 per share. Goode’s dividend is expected to grow at a constant rate of 7.00%. What was the last dividend, D0? Please show all work and formulas used.

a. $0.95

b. $1.05

c. $1.16

d. $1.27
Answer #1

Ans b. $1.05

D1 = Current Dividend
Ke = Cost of Equity
g = growth rate
P0 = D1 / (Ke - g)
25 = D1 / (11.5% - 7%)
D1 = 25 * 4.5%
D1 = 1.125
D1 = D0 * (1 + g)
1.125 = D0 * ( 1 + 7%)
D0 = 1.125 / 1.07
D0 = 1.05